If the answer to a subtraction problem is 24,list five possible number combinations that were in the problem.
Answer by fcabanski(1391)

x-y = 24.


48 and 24: 48 - 24 = 24


100 and 76: 100 - 76 = 24


-10 and -34: -10 - (-34) = 24


34, 6 and 4: 34 - 6 - 4 = 24


24.5 and .5: 24.5 - .5 = 24